2). What is the hydronium ion concentration in an aqueous solution with a hydroxide ion concentration of 0.01 Mat 25 °C? O A)What is the hydronium ion concentration in an aqueous solution with a hydroxide ion concentration of 0.01 M at 25 °C?
Answer #1

2) Answer is E)

[OH-] = 0.01 = 10-2 M

[H3O+] = Kw/[OH-] = 1×10-14/10-2 = 1.0×10-12 M

Where Kw is ionic product of water = 1×10-14